OpenStudy (anonymous):

I am trying to help my son. A canoe travels 4 miles per hour in still water. It travels 6 miles up stream and back again in 4 hours. what is the speed of the current? He got an answer of 5.2 is that correct?

OpenStudy (mathmate):

It cannot be correct because 5.2 is greater than 4, so he will never go upstream. Let s= speed of current, then \[\frac{6}{4-s}+\frac{6}{4+s}=4\] solve for s. (hint: reject negative root)

OpenStudy (anonymous):

but it is less than 8 (4 up and 4 back again)

OpenStudy (anonymous):

Im getting him......

OpenStudy (mathmate):

Well, if you (are trying to) go upstream and can row at 4 mph. Current is rushing at you at 5.2 mph, you are going back at 1.2 mph, so you'll never get upstream. This has nothing to do with the 4 hours it takes to go up- and downstream.

OpenStudy (anonymous):

if I scan his work, can you tell us where he went wrong?

OpenStudy (anonymous):

it is not posting, it just hangs up

OpenStudy (anonymous):

I scanned it as a different file type and it worked

OpenStudy (mathmate):

He works well. Everything is correct, except he made a minor mistake at the very beginning. Knowing that he has to row upstream, we would expect that the current r should be less than 4. The first line should read 6(4-r)+6(4+r)=4(4-r)(4+r) and everything will turn out correctly.

OpenStudy (anonymous):

THANK YOU!

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

we ended up stuck again at 0=16-4r²

OpenStudy (mathmate):

Divide by 4 on both sides: \[0=4-r^2\] add r^2 to both sides and take square-root \[r^2=4\] Let him try finishing the rest. (hint: reject negative root)

OpenStudy (anonymous):

r=2

OpenStudy (anonymous):

?

OpenStudy (mathmate):

The answer is correct. Check the time it takes to do the return trip: 6/(4-2)+6/(4+2)=3+1=4 hours...ok

OpenStudy (anonymous):

are you available to tell me if he got the right answer on the next problem?

OpenStudy (anonymous):

A boat travels 6 miles per hour in still water. It travels 8 miles upstream and back again in 8 hours. what is the speed of the current? r=4.9

OpenStudy (mathmate):

Correct to 1 decimal place. To be exact, it is \[2\sqrt{6}=4.899...\]

OpenStudy (anonymous):

yes! he rounded up. Again thank you. nothing like having a bunch of math I cant do that must be done by tomorrow in order to pass the class! 2 down, 15 to go